OpenStudy (anonymous):

which statement is true? all irrational numbers are also rational numbers irrational numbers cannot be classified as rational numbers

OpenStudy (anonymous):

Rational and irrational numbers are opposites. You were given a link to wikipedia earlier, you really should read it.

OpenStudy (anonymous):

what the hell I blocked you any ways

OpenStudy (anonymous):

i think this is a reading comprehension problem math is weird, but it is not that weird how could an "irrational" number also be "rational"?

OpenStudy (anonymous):

i have no idea right it cant be both

OpenStudy (anonymous):

You blocked me? No problem. I'll try to remember not to answer anymore questions from you.

OpenStudy (anonymous):

thank god you don't even answer them anyways you just send me somewhere else

OpenStudy (anonymous):

Well, I know that "all irrational numbers are also rational numbers" Is false. A counterexample to this would be "The Euler Mascheroni constant" defined as: \[\int^\infty_1{\left(\frac{1}{\lfloor{x}\rfloor}-\frac{1}{x}\right)}\phantom{0}dx\]

OpenStudy (anonymous):

And since "If either a or b is true and a is false, therefore b must be true" I guess, irrational numbers cannot be classified as rational numbers

OpenStudy (debbieg):

Holy cr@p, can't we just use \(\sqrt{2}\) as a counterexample?? LOL!

OpenStudy (anonymous):

thank you so much so irrational numbers cannot be classified as rational numbers

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks to everyone that helped me ..

OpenStudy (anonymous):

that is why they are called "irrational" i.e. "not rational"

OpenStudy (anonymous):

brain fart right. i read it and reread that stupid question and walked away i got it . with your guys help thanks

OpenStudy (anonymous):

Haha, well in a sense, \(\sqrt{2}=\frac{2}{\sqrt{2}}\) I considered using \(e\) but then I thought that \[e=\frac{1}{0!}+\frac{1}{1!}+\frac{1}{2!}+\frac{1}{3!}+\frac{1}{4!}...\] And realized I wanted a more, less rational strong counterexample lol

OpenStudy (anonymous):

lol